oregonduck (oregonduck):

Pre- Algebra Help @misty1212 could you help?

oregonduck (oregonduck):

A closed container has 3.06 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10−24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 0.47 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10−24) = 4.73 ⋅ 10−1 0.51 grams, because (3.06 ⋅ 1.67) ⋅(1023 ⋅ 10-24) = 5.1102 ⋅ 10-1 4.73 grams, because (3.06 + 1.67) ⋅ (1023 ⋅ 10−24) = 4.73 5.11 grams, because (3.06 ⋅ 1.67) ⋅ (1023 ⋅ 10−24) = 5.1102

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i can make a guess

oregonduck (oregonduck):

the number after the 10's are exponts fyi

OpenStudy (misty1212):

first off it is probably \[3.06\times 10^{-23}\] right? and \[1.67\times 10^{-24}\]

oregonduck (oregonduck):

correct

OpenStudy (misty1212):

on no first one is \[3.06\times 10^{23}\] no minus sign

oregonduck (oregonduck):

yea lol forgot

OpenStudy (misty1212):

so you have to multiply \[3.06\times 10^{23}\times 1.67\times 10^{-24}\] which is the same as \[\huge (3.06\times 1.67)\times 10^{23-24}\]

OpenStudy (misty1212):

need a calculator for the first part for sure

oregonduck (oregonduck):

( 5.1102

oregonduck (oregonduck):

So then D :D

OpenStudy (misty1212):

well maybe \[5.1102\times 10^{-1}\]

oregonduck (oregonduck):

5.11 grams, because (3.06 ⋅ 1.67) ⋅ (1023 ⋅ 10−24) = 5.1102

OpenStudy (misty1212):

so no, not D

OpenStudy (misty1212):

move the decimal over one to the left, gives \[0.51102\]

oregonduck (oregonduck):

oh ok then B because of the 10^-1

oregonduck (oregonduck):

\[0.51 becasue 5.1102 * 10^-1\]